OpenStudy (anonymous):

2^x+3=5^x If someone could take me step by step in this problem????

OpenStudy (anonymous):

okay, so in this problem, you have to be familiar with the log function (which i assume you are)

OpenStudy (anonymous):

A little, thats what we are learning now

OpenStudy (anonymous):

I know that I have to natural (ln) each side..

OpenStudy (anonymous):

I came up with an answer of 2.269 I just know its wrong...

OpenStudy (anonymous):

if someone is good with vectors here, so plz give some thought to my question as well.

OpenStudy (anonymous):

sorry, logs are so far back in my math knowledge. im trying to remember the rules

OpenStudy (nikvist):

\[2^x+3=5^x\] or \[2^{x+3}=5^x\]

OpenStudy (anonymous):

Here is my work... maybe someone can check it to see where I went wrong. 2^(x+3)=5^x Then I divided each side by x ln 2 and it left me with: x+3/x=ln5/ln2

OpenStudy (anonymous):

The second one...

OpenStudy (anonymous):

ohhh i remember how to do it. i thought it was 2^x + 3 instead of 2^(x+3)

OpenStudy (anonymous):

I dont know how to write equations on here.... so sorry

OpenStudy (anonymous):

its cool. alright, so log both sides to start. that will get you ln(2^(x+3)) = ln(5^x)

OpenStudy (anonymous):

then with the logarithm rules, you can bring the exponent of both sides in front. this gives you: (x+3)ln(2) = xln(5)

OpenStudy (anonymous):

then you switch things around to get (x+3)/x = ln(5)/ln(2)

OpenStudy (anonymous):

Ok, I got that. So far, so good

OpenStudy (anonymous):

then you can split the fraction on the left up to make two different ones. It becomes (x/x) + (3/x)

OpenStudy (anonymous):

or, in other words, (3/x) + 1

OpenStudy (anonymous):

OK... right here is where my teacher told us to croos multiply?

OpenStudy (anonymous):

cross sorry. So it gave me x+3/x=2.322/1

OpenStudy (anonymous):

does your teacher want you to get an exact answer? or are decimals okay

OpenStudy (anonymous):

Decimals to the thousandth

OpenStudy (anonymous):

okay, well i would highly recommend solving for x in exact form first, and then getting the decimal. it's much more pure math. if you have (x/3) + 1 = ln(5)/ln(2) it's very easy to get it to exact form

OpenStudy (anonymous):

sorry, (3/x)

OpenStudy (anonymous):

just move the 1 over to the right side

OpenStudy (anonymous):

What one? where are you getting the one from?

OpenStudy (anonymous):

then it will be 3/((ln(5)/ln(2))+1)

OpenStudy (anonymous):

from the left side. where it says (3/x) + 1 = ln(5)/ln(2)

OpenStudy (anonymous):

it becomes 3/x = ln(5)/ln(2) -1

OpenStudy (anonymous):

which becomes 3/((ln(5)/ln(2))-1)

OpenStudy (anonymous):

^^ that is equal to x

OpenStudy (anonymous):

is the answer 1.689?

OpenStudy (anonymous):

i believe its 2.269. make sure you plug it in correctly on your calculator.

OpenStudy (anonymous):

Thats what I had for my answer to begin with?

OpenStudy (anonymous):

So I am right?

OpenStudy (anonymous):

haha, yes

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (anonymous):

no problem